If a consumer daily income rises from Rs. 300 to Rs. 350, his purchase of a good X increases 

from 25 units to 35 units, find income elasticity of demand for X.


1
Expert's answer
2022-02-21T13:32:47-0500

To determine the elasticity of demand for the needed commodities, we shall apply the straightforward formula. We already have the change in income amount and the change in demand, so we'll compute the needed entity using the direct values in the formula.


If a customer's daily income grows from Rs. 300 to Rs. 350, the purchase of a good X increases from 25 to 35 units per day.


Assume that the demand is represented by the letter "Q."

Similarly, the revenue will be denoted by the letter "M".


Let's say "Q1" is "25" and "Q2" is "35" .

As a result, the change in demand is determined by:


Δ"Q" ="Q""2"-"Q1" ⇒Δ"Q" "=10"


Assume "M1""=300" and "M2=350" in the same way.

As a result, the change in income is determined by:


Δ"M" ="M2""M1" ⇒Δ"M=50"

The income elasticity is now calculated using the following formula:


"=\frac{ΔQ}{ΔM}\times\frac{M2+M1}{Q2+Q1}"


By substituting the aforementioned values into the formula, we get:

"=\frac{10}{50}\times\frac{650}{60}"

"=2.16666667"